Grade upto college level Physical Chemistry

Sodium metal crystallizes in body centred cubic lattice with the cell edge, a = 4.29Å. What is the redius of sodium atom?

Shane Macguire
12 Years agoGrade upto college level

3 Answers

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
For bcc lattice, (radius), r = √3a/4
Solution
∴ r = √3 *4.29 Å/4 = 1.73 * 4.29 Å/4 = 1.86

Thanks
Deepak patra
askIITians Faculty
Aparna
7 Years ago
1.73× 42.29= 73.1617 .         ✓3 = 1.73
   73.1617/4 = 18.3 ( considered only one digit after decimal) 
 
Hope it helps.  .......
Kushagra Madhukar
6 Years ago
Dear student,
Please find the answer to your problem
 
For bcc lattice, (radius), r = √3a/4
∴ r = √3 *4.29 Å/4 = 1.73 * 4.29 Å/4 = 1.86
 
Hope it helps.
Thanks and regards,
Kushagra